if A(vector) = 3i + 4j and B(vector) = 7i + 24j , then what wiil be the vector that has same magnitude as A(vector) and direction parallel to B(vector).

 if  A(vector) = 3i + 4j    and     B(vector) = 7i + 24j ,


then what wiil be the vector that has same magnitude as A(vector) and direction parallel to B(vector).

Grade:11

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Neeru,

 Ans:- The vector must be=p(7i+24j) where p is a cons tant

Hence p√(24²+7²)=√3²+4² ( according to the given condition)

or p25=5

or p=1/5

Hence the vector is=(7i+4j)/5 (Ans)

 


Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Neeru !!!

 


Regards,

Askiitians Experts

SOUMYAJIT IIT-KHARAGPUR